# Operator Spreading and the Absence of Many-Body Localization

### v2

@@ Abstract
We consider the spreading of a local operator $A$ in Euclidean time in one-dimensional many-body systems with Hamiltonian $H$ by calculating the $k$-fold commutator $[H,[H,[...,[H,A]]]]$. We derive general bounds for the operator norm of this commutator in free and interacting fermionic systems with and without disorder. We show, in particular, that in a localized system the norm does grow at most exponentially and that the contributions of operators to the total norm are exponentially suppressed with their length. We support our general results by considering one specific example, the XXZ chain with random magnetic fields. We solve the operator spreading in the XX case without disorder exactly. For the Anderson and Aubry-André models we provide strict upper bounds. We support our results by symbolic calculations of the commutator up to high orders. For the XXX case with random magnetic fields, these symbolic calculations show a growth of the operator norm faster than exponential and consistent with the general bound for a non-localized system. Also, there is no exponential decay of the contribution of operators as function of their length. We conclude that there is no indication for a many-body localization transition. Finally, we also discuss the differences between the interacting and non-interacting cases when trying to perturbatively transform the microscopic to an effective Hamiltonian of local conserved charges by consecutive Schrieffer-Wolff transformations. We find that such an approach is not well-defined in the interacting case because the transformation generates $\sim 4^{\ell}$ terms connecting sites a distance $\ell$ apart which can overwhelm the exponential decay with $\ell$ of the amplitude of each individual term.
@@ I Introduction
Here we want to propose and follow a new avenue by studying the properties of operators instead of wave functions. Our motivation is the following observation: In an Anderson localized model, the projectors onto the eigenstates $P_{n}=|E_{n}\rangle\langle E_{n}|$ are conserved quasi-local operators, i.e., they are centered at a site in the lattice and decay exponentially away from this site. These operators are also often called local integrals of motion (LIOMs). Without loss of generality, we can order the operators such that $P_{n}$ is an operator localized around lattice site $n$. To be more precise, we can use the fermionic language in which the Anderson Hamiltonian and any unitary transformation of this Hamiltonian is bilinear. We can then write $P_{n}=\sum_{ij}a^{n}_{ij}c^{\dagger}_{i}c_{j}$ where $c_{i}^{(\dagger)}$ is a fermionic annihilation (creation) operator at site $i$ and $|a^{n}_{ij}|\sim\exp(|i-n|)\exp(|j-n|)$ are amplitudes. In particular, the contribution of bilinear operators to $P_{n}$ is exponentially suppressed with the distance they cover.
There is a unitary transformation $U$ which diagonalizes the microscopic, non-interacting Anderson Hamiltonian $H$, bringing it into the form $\tilde{H}=UHU^{-1}=\sum_{n}E_{n}P_{n}$ where $P_{n}$ is quasi-local and conserved, $[H,P_{n}]=0$. Importantly, this unitary transformation itself is quasi-local and will always map local microscopic operators $O_{i}^{\alpha}$ onto quasi-local operators, $\tilde{O}^{\alpha}_{i}=UO^{\alpha}_{i}U^{-1}$. As we will show, this has important implications for the spreading of a local operator in Euclidean time. We can write
$O_{i}^{\alpha}(\tau)$ $=$ $\text{e}^{\tau H}O^{\alpha}_{i}\text{e}^{-\tau H}$
$=$ $\sum_{k=0}^{\infty}\underbrace{[H,[H,[...,[H,O_{i}^{\alpha}]]]]}_{\textrm{k-times}}\frac{\tau^{k}}{k!}.$
I.e., the spreading of the operator is encoded in the $k$-fold commutator $[H,O_{i}^{\alpha}]^{(k)}$ with the Hamiltonian. It is important to note that the Frobenius norm of this commutator is not affected by the unitary transformation $||[H,O_{i}^{\alpha}]^{(k)}||_{2}=||[\tilde{H},\tilde{O}_{i}^{\alpha}]^{(k)}||_{2}$. For the latter we expect that the norm of operators in the commutator which cover a distance of $l$ sites is exponentially suppressed with $l$ at fixed order $k$. As we will show, this implies that the total norm of all terms in the commutator can only grow exponentially with $k$.
@@
This has to be contrasted with the general, non-localized case. In one dimension, there is no phase transition in a clean system at finite temperatures. Thus, the time evolution is always analytic and Eq. (I) implies that the norm of the $k$-fold commutator $[H,O_{i}^{\alpha}]^{(k)}$ has to grow slower than factorially. In Ref. [30], it has been shown that this is indeed the case, however, the derived bound for the total norm grows much faster than exponential and almost factorially. It has been hypothesized that any generic quantum chaotic system will asymptotically saturate this bound [30, 31]. Thus, we have a sharp contrast between the growth of the operator norm of the $k$-fold commutator of a local operator with the Hamiltonian between a localized system, where it can grow at most exponentially, and a non-localized system, where an almost factorial growth is expected in general. This qualitative difference in the operator norm growth is what we will investigate here. To avoid misunderstandings, we want to stress already here that what we mean by a local operator and what we will investigate in the following is an operator $O^{\alpha}_{j}$ which resides on lattice site $j$ in the microscopic model. In contrast, operators of the form $O^{\alpha}=\sum_{j}O^{\alpha}_{j}$ are also often called local conserved charges if $[H,O^{\alpha}]=0$ and are of particular importance in integrable models. Operators of the type $O^{\alpha}$, which are sums of local densities but which extend over the entire lattice, are also the type of operators whose growth is investigated in Refs. [31, 32]. The latter type of operators are not of interest for a study of localization and we will not consider them here.
@@
In addition to analytical results, we will also present exact results from symbolic calculations of commutators and norms. To be able to consider high orders $k$ in the commutator, we have written code in Julia with Nemo [33] which is able to handle symbolic computations with up to $\sim 200$ million terms, limited only by the available memory. We note that we are able to consider system sizes for nearest-neighbor models which are about twice as large as the ones typically considered in exact diagonalizations.
@@
Our paper is organized as follows: In Sec. II, we discuss general bounds for the growth of the matrix norm $||[H,A]^{(k)}||$ for a local operator $A$ and a nearest-neighbor Hamiltonian $H$. We review the result in the general interacting case without localization [30] before discussing the free-fermion case and the case of localization. In Sec. III, we then consider a specific model, the $XXZ$ spin chain. We start with the free-fermion (XX) model without disorder before discussing the Anderson case of quenched disorder. We also consider the Aubry-André model which has a periodic potential incommensurate with the lattice and shows a phase transition at a finite disorder strength which, as we will demonstrate, is visible in the norm growth. We then turn to the interacting Heisenberg (XXX) model and show that results obtained by symbolic computations are inconsistent with localization for all studied disorder strengths. In Sec. IV, we consider Schrieffer-Wolff transformations which can be used to perturbatively transform the Hamiltonian into a basis of quasi-local conserved operators. We show that such a perturbative construction succeeds in the non-interacting case while it will typically fail once interactions are included. The final section contains a brief summary of our main results and a discussion of the remaining open issues.
@@ II Bounds on operator norm growth
We consider two entry-wise matrix norms for a matrix $A$ with elements $a_{i,j}$:
@@ II.1 General Case
There are $2^{k}$ terms of this kind. If we use again the sub-multiplicativity property then we obtain
$||[H,A]^{(k)}||$$\leq$ $\sum_{\{I_{1},\cdots,I_{k}\}}\sum_{\{a_{I}\}}\underbrace{||h||\cdots||h||}_{\text{$k$-many}}\,||A||$ (II.10)
@@
where we have used that $\sum_{\{a_{I}\}}=C^{k}$ and $||h^{a}_{I}||=J^{a}_{I}\leq J=\text{max}_{I,a}\{J^{a}_{I}\}$.
@@
$s(k)\equiv||[H,A]^{(k)}||\leq||A||(2CJ)^{k}B_{k}(2)$ (II.12)
where norm|$\dots$||| | … | | can be either the $1$-norm or the $2$-norm defined earlier. We note that this bound is also valid for any sub-multiplicative norm.
@@
We can also ask the question how many distinct terms are possible in the commutator $[H,A]^{(k)}$. Note that this question is different from the question how many connected clusters exist which we addressed above because connected clusters formed in different ways can result in the same operator. Let us assume that our system has a basis of $f-1$ local operators plus the local identity $\mathbb{I}_{j}$. At order $k$, clusters have at most length $k+1$ and, if the operator $A$ was placed on site $j=0$, extend from sites $-k,-k+1,\dots,0$ to the right. I.e., there are $k+1$ clusters of length $k+1$ with each shifted by one lattice site to the right compared to the previous one. For the cluster starting at $-k$, we can put any of the $f-1$ local operators or the identity on each of the $k+1$ sites leading to $f^{k+1}$ possible operators. For the cluster starting at site $k-1$, we have to put one of the $f-1$ local operators at the right most site, $j=1$, to obtain operators distinct from the ones already constructed. Thus, this cluster will add $(f-1)f^{k}$ possible operators. The same construction also works for the remaining $k-1$ clusters. We can therefore bound the total number of possible distinct operators occuring in the commutator $[H,A]^{(k)}$ by
@@
For a generic system, we thus expect that the number of distinct terms in the commutator grows exponentially. If such a system does not have local operators which are conserved, then the hypothesis is that the bound on the total norm (II.12) is asymptotically tight [30]. We note again that by local we mean operators acting on one or a few neighboring sites in the microscopic model, not operators which are sums of local densities. Next, we consider non-interacting systems where stricter, exponential bounds on the norm can be derived.
@@ II.2 Free Fermions
In a free-fermion system, the number of distinct operators in the commutator $[H,A]^{(k)}$ thus grows quadratically with the order of the commutator $k$.
@@ II.3 Localization
$\sum_{j=1}^{\xi_{\rm loc}}2^{j}S(k,j)\sim\sum_{j=1}^{\xi_{\rm loc}}\frac{2^{j}j^{k}}{k!}<\frac{2^{\xi_{\rm loc}}}{(\xi_{\rm loc}-1)!}\xi_{\rm loc}^{k}$ (II.20)
@@
We note that both the free-fermion case and the localized case show an exponential growth of the operator norm. Thus, the total operator norm is insufficient to distinguish a non-localized free-fermion phase from a localized interacting or non-interacting phase. What does distinguish the two phases though is how different terms in the commutator contribute to the norm. If we denote by $[H,A]^{(k)}_{l}$ the terms in the commutator of order $k$ which have support on $l$ lattice sites, then we expect that in the localized case
$s_{l}(k)\equiv||[H,A]^{(k)}_{l}||\sim\text{e}^{k}\text{e}^{-l}\,.$ (II.22)
@@
To summarize, we hypothesize that the bound (II.12) is asymptotically tight for an interacting, non-localized system which thus should have an operator norm $s(k)$ which grows faster than exponential with the order of the commutator $k$. In an interacting localized phase—the putative many-body localized phase—the norm growth, as we have proven, can instead only be at most exponential. For a free-fermion model, the norm growth is also always at most exponential whether it is in a localized phase or not. In the latter case, we can distinguish a localized from a non-localized phase by studying the contributions $s_{l}(k)$ from terms which are supported on $l$ lattice sites to the total norm of the commutator $s(k)$.

### v3

# Operator Growth in Disordered Spin Chains: Indications for the Absence of Many-Body Localization

@@ Abstract
We consider the growth of a local operator $A$ in one-dimensional many-body systems with Hamiltonian $H$ by calculating the $k$-fold commutator $[H,[H,[...,[H,A]]]]$. We derive general bounds for the operator norm of this commutator in free and interacting fermionic systems with and without disorder thus directly connecting the operator growth hypothesis with questions of localization. We show, in particular, that in a localized system the norm does grow at most exponentially and that the contributions of operators to the total norm are exponentially suppressed with their length. We support our general results by considering one specific example, the XXZ chain with random magnetic fields. We solve the operator spreading in the XX case without disorder exactly. For the Anderson and Aubry-André models we provide strict upper bounds. We support our results by symbolic calculations of the commutator up to high orders. For the XXX case with random magnetic fields, these symbolic calculations show a growth of the operator norm faster than exponential and consistent with the general bound for an ergodic system. Also, there is no exponential decay of the contribution of operators as function of their length. We conclude that there is no indication for a many-body localization transition. Finally, we also discuss the differences between the interacting and non-interacting cases when trying to perturbatively transform the microscopic to an effective Hamiltonian of local conserved charges by consecutive Schrieffer-Wolff transformations. We find that such an approach is not well-defined in the interacting case because the transformation generates $\sim 4^{\ell}$ terms connecting sites a distance $\ell$ apart which can overwhelm the exponential decay with $\ell$ of the amplitude of each individual term.
@@ I Introduction
Here we want to propose and follow a new avenue by studying the properties of operators instead of wave functions. Our motivation is the following observation: In an Anderson localized model, the projectors onto the eigenstates $P_{n}=|E_{n}\rangle\langle E_{n}|$ are conserved quasi-local operators, i.e., they are centered at a site in the lattice and decay exponentially away from this site. These operators are also often called local integrals of motion (LIOMs). Without loss of generality, we can order the operators such that $P_{n}$ is an operator localized around lattice site $n$. To be more precise, we can use the fermionic language in which the Anderson Hamiltonian and any unitary transformation of this Hamiltonian are bilinear. We can then write $P_{n}=\sum_{ij}a^{n}_{ij}c^{\dagger}_{i}c_{j}$ where $c_{i}^{(\dagger)}$ is a fermionic annihilation (creation) operator at site $i$ and $|a^{n}_{ij}|\sim\exp(|i-n|)\exp(|j-n|)$ are amplitudes. In particular, the contribution of bilinear operators to $P_{n}$ is exponentially suppressed with the distance they cover.
There is a unitary transformation $U$ which diagonalizes the microscopic, non-interacting Anderson Hamiltonian $H$, bringing it into the form $\tilde{H}=UHU^{-1}=\sum_{n}E_{n}P_{n}$ where $P_{n}$ is quasi-local and conserved, $[H,P_{n}]=0$. Importantly, this unitary transformation itself is quasi-local and will always map local microscopic operators $A_{i}$ onto quasi-local operators, $\tilde{A}_{i}=UA_{i}U^{-1}$. As we will show, this has important implications for the spreading of a local operator in Euclidean time. We can write
$A_{i}(\tau)$ $=$ $\text{e}^{\tau H}A_{i}\text{e}^{-\tau H}$
$=$ $\sum_{k=0}^{\infty}\underbrace{[H,[H,[...,[H,A_{i}]]]]}_{\textrm{k-times}}\frac{\tau^{k}}{k!}.$
I.e., the spreading of the operator is encoded in the $k$-fold commutator $[H,A_{i}]^{(k)}$ with the Hamiltonian. It is important to note that the Frobenius norm of this commutator is not affected by the unitary transformation $||[H,A_{i}]^{(k)}||_{2}=||[\tilde{H},\tilde{A}_{i}]^{(k)}||_{2}$. For the latter we expect that the norm of operators in the commutator which cover a distance of $l$ sites is exponentially suppressed with $l$ at fixed order $k$. As we will show, this implies that the total norm of all terms in the commutator can only grow exponentially with $k$.
@@
This has to be contrasted with the general, ergodic case. In one dimension, there is no phase transition in a clean system at finite temperatures. Thus, the time evolution is always analytic and Eq. (I) implies that the norm of the $k$-fold commutator $[H,A_{i}]^{(k)}$ has to grow slower than factorially. In Ref. [30], it has been shown that this is indeed the case, however, the derived bound for the total norm grows much faster than exponential and almost factorially, $||[H,A_{i}]^{(k)}||_{2}\sim(k/\ln k)^{k}$. The operator growth hypothesis states that any generic quantum chaotic system will asymptotically saturate this bound [30, 31]. Thus, we have a sharp contrast between the growth of the operator norm of the $k$-fold commutator of a local operator with the Hamiltonian between a localized, non-ergodic system, where it can grow at most exponentially, and a non-localized system, where an almost factorial growth is expected in general. This qualitative difference in the operator norm growth is what we will investigate here. To avoid misunderstandings, we want to stress already here that what we mean by a local operator and what we will investigate in the following is an operator $A_{i}$ which resides on lattice site $i$ in the microscopic model. In contrast, operators of the form $A=\sum_{i}A_{i}$ are also often called local conserved charges if $[H,A]=0$ and are of particular importance in integrable models. Operators of the type $A$, which are sums of local densities but which extend over the entire lattice, are also the type of operators whose growth has been investigated in Refs. [31, 32]. The latter type of operators are not of interest for a study of localization and we will not consider them here.
We note that in contrast to real-time evolution, the Euclidean time evolution of operators can be directly connected to questions of quantum chaos and the eigenstate thermalization hypothesis (ETH) [30, 31]. First, the Frobenius norm of a local operator $A$ under Euclidean time evolution in the energy eigenbasis for a finite system of size $L$ is given by
[…]
For a system satisfying ETH, most matrix elements $A_{ij}$ are expected to be non-zero even for extensive energy differences $|E_{i}-E_{j}|$. This implies that $||A(\tau)||_{2}^{2}\sim\exp(J\tau L)$ where $J$ is the local energy scale. If a system is non-interacting or has local conservation laws on the other hand, and thus does not satisfy ETH, matrix elements will be zero for extensive energy differences and we thus expect $||A(\tau)||_{2}^{2}\sim\exp(J\tau)$. By deriving strict bounds directly for the Frobenius norm $||A(\tau)||_{2}$, one can use Eq. (I.3) to make precise statements about the scaling of off-diagonal matrix elements $A_{ij}$ as a function of $|E_{i}-E_{j}|$. For the generic, chaotic case this scaling has been considered in Ref. [30] and is consistent with numerical investigations [33, 34].
Another quantity which can be obtained from considering the Euclidean time evolution of local operators are Lieb-Robinson bounds. A clean, one-dimensional system with short-range interactions shows a linear spreading of information under real-time evolution independent of whether the system is chaotic or not [35, 36]. In contrast, under Euclidean time evolution information spreads exponentially fast, $\ell\sim\text{e}^{\tau}$, over a distance $\ell$ in a 1D chaotic system [30] while it only spreads ballistically in a non-interacting system or a system with local conservation laws. It has also been speculated that the time for the spreading of information in Euclidean time in a 1D chaotic system, $\tau\sim\ln(\ell)$, is indicative of the Thouless time in a system without conservation laws (for example a Floquet system), i.e., the time after which the system can be described by a random matrix. Furthermore, it is also possible to connect the characteristic time of Euclidean operator growth with the delocalization of this operator in Krylov space [31]. Overall, there is strong evidence that in a quantum chaotic system, an operator grows asymptotically at its maximally allowed rate and that a parametrically slower growth is related to ergodicity breaking.
Here we want to expand the study of operator growth to disordered systems. There is, in particular, a direct conflict between the operator growth hypothesis which suggests that an almost factorial growth is indicative of ergodicity and systems such as the random transverse Ising model which has been suggested to have a many-body localized phase but does always show almost factorial operator growth [37]. While we will consider Lieb-Robinson bounds and the decay of off-diagonal matrix elements in App. B and App. C, it is important to note that these quantities are obtained by taking sums over $k$-fold commutators $[H,A_{i}]^{(k)}$ and thus some of the information, in particular about the spatial support of such commutators which is relevant for the question of localization, is lost. Here we follow a novel path by directly obtaining analytical results for the asymptotic scaling of $k$-fold commutators resolved by the extent of their spatial support. Our results resolve the above mentioned conflict and suggest that the operator growth hypothesis is valid and that many-body systems, even for strong disorder, do not localize but rather stay ergodic.
@@
In addition to these analytical results, we will also present exact results from symbolic calculations of commutators and norms. To be able to consider high orders $k$ in the commutator, we have written code in Julia with Nemo [38] which is able to handle symbolic computations with up to $\sim 200$ million terms, limited only by the available memory. Nemo also provides arbitrary precision complex ball arithmetic which tracks error bounds rigorously. We typically use $200$ or $400$ digit precision for the amplitudes of the operators. While the absolute errors for large commutator orders can become large, we keep the relative errors always below $10^{-50}$ such that error bars are not visible in many of the plots presented in the following. We note that by using symbolic calculations we are able to consider system sizes for nearest-neighbor models which are about twice as large as the ones typically considered in exact diagonalizations.
@@
Our paper is organized as follows: In Sec. II, we discuss general bounds for the growth of the norm $||[H,A]^{(k)}||$ for a local operator $A$ and a nearest-neighbor Hamiltonian $H$. We review the result in the general interacting case without localization [30] before discussing the free-fermion case and the case of localization. In Sec. III, we then consider a specific model, the $XXZ$ spin chain. We start with the free-fermion (XX) model without disorder before discussing the Anderson case of quenched disorder. We also consider the Aubry-André model which has a periodic potential incommensurate with the lattice and shows a phase transition at a finite disorder strength which, as we will demonstrate, is clearly visible in the norm growth. We then turn to the interacting Heisenberg (XXX) model and show that results obtained by symbolic computations are inconsistent with localization for all studied disorder strengths. In Sec. IV, we consider Schrieffer-Wolff transformations which can be used to perturbatively transform the Hamiltonian into a basis of quasi-local conserved operators. We show that such a perturbative construction succeeds in the non-interacting case while it will typically fail once interactions are included. The final section contains a summary of our main results and a discussion of the remaining open issues.
@@ II Bounds on operator norm growth
We consider two entry-wise matrix norms for a matrix $A$ with elements $A_{ij}$:
@@ II.1 General Case
There are $2^{k}$ terms of this kind and we therefore find
$||[H,A]^{(k)}||$$\leq$ $2^{k}\sum_{\{I_{1},\cdots,I_{k}\}}\sum_{\{a_{I}\}}\underbrace{||h||\cdots||h||}_{\text{$k$-many}}\,||A||$ (II.10)
@@
where we have used that $||h^{a}_{I}||=J^{a}_{I}\leq J=\text{max}_{I,a}\{J^{a}_{I}\}$ and $\sum_{\{a_{I}\}}=C^{k}$.
@@
$s^{(p)}(k)\equiv||[H,A]^{(k)}||_{p}\leq||A||_{p}(2CJ)^{k}B_{k}(2)$ (II.12)
where $p=1,2$ is the $1$-norm or the $2$-norm defined earlier. We note that this bound is also valid for any sub-multiplicative norm.
@@
We can also ask the question how many distinct terms are possible in the commutator $[H,A]^{(k)}$. Note that this question is different from the question of how many connected clusters exist which we addressed above because connected clusters formed in different ways can result in the same operator. Let us assume that our system has a basis of $f-1$ local operators plus the local identity $\mathbb{I}_{j}$. At order $k$, clusters have at most length $k+1$ and, if the operator $A$ was placed on site $j=0$, extend from sites $-k,-k+1,\dots,0$ to the right. I.e., there are $k+1$ clusters of length $k+1$ with each shifted by one lattice site to the right compared to the previous one. For the cluster starting at $-k$, we can put any of the $f-1$ local operators or the identity on each of the $k+1$ sites leading to $f^{k+1}$ possible operators. For the cluster starting at site $-k+1$, we have to put one of the $f-1$ local operators at the right most site, $j=1$, to obtain operators distinct from the ones already constructed. Thus, this cluster will add $(f-1)f^{k}$ possible operators. The same construction also works for the remaining $k-1$ clusters. We can therefore bound the total number of possible distinct operators occuring in the commutator $[H,A]^{(k)}$ by
@@
For a generic system, we thus expect that the number of distinct terms $n(k)$ in the commutator grows exponentially. As we will discuss in more detail below, the terms occurring in the commutator can be represented as nodes in a graph and $n(k)$ is directly related to the complexity of this graph.
To summarize, if a system does not have local operators which are conserved, then the hypothesis is that the bound on the total norm (II.12) is asymptotically tight [30]. Importantly, a norm growth $s(k)\sim(k/\ln k)^{k}$ is not possible in regular graphs or Cayley trees where the norm cannot grow faster than exponential. Many-body systems and the question of localization in such systems are therefore qualitatively different from studies of localization on regular graphs [40, 41]. Next, we consider non-interacting systems which are one example of systems corresponding to regular graphs where stricter, exponential bounds on the norm can be derived.
@@ II.2 Free Fermions
In a free-fermion system, the number of distinct operators in the commutator $[H,A]^{(k)}$ thus grows quadratically with the order of the commutator $k$. As we will show below when considering the example of the XX chain, the quadratic growth of the number of distinct operators is directly related to a regular structure of the corresponding graph with a constant number of nodes and edges at each level.
@@ II.3 Localization
$\sum_{j=1}^{\xi_{\rm loc}}2^{j}S(k,j)\sim\sum_{j=1}^{\xi_{\rm loc}}\frac{2^{j}j^{k}}{k!}<\frac{2^{\xi_{\rm loc}}}{(\lceil\xi_{\rm loc}\rceil-1)!}\xi_{\rm loc}^{k}$ (II.20)
@@
We note that both the free-fermion case and the localized case show an exponential growth of the operator norm. Thus, the total operator norm is insufficient to distinguish a non-localized free-fermion phase from a localized interacting or non-interacting phase. The total norm, however, does tell us whether or not a system is ergodic. What does distinguish non-ergodic non-localized from localized phases though is how different terms in the commutator contribute to the norm. If we denote by $[H,A]^{(k)}_{l}$ the terms in the commutator of order $k$ which have support on $l$ lattice sites, then we expect that in the localized case
$s^{(p)}_{l}(k)\equiv||[H,A]^{(k)}_{l}||_{p}\sim\text{e}^{k}\text{e}^{-l}\,.$ (II.22)
@@
To summarize, the operator growth hypothesis is that the bound (II.12) is asymptotically tight for an ergodic system which thus should have an operator norm $s(k)$ which grows faster than exponential with the order of the commutator $k$. In an interacting localized phase—the putative many-body localized phase—the norm growth, as we have proven, can instead only be at most exponential. For a free-fermion model, the norm growth is also always at most exponential whether it is in a localized phase or not. In the latter case, we can distinguish a localized from a non-localized phase by studying the contributions $s_{l}(k)$ from terms which are supported on $l$ lattice sites to the total norm of the commutator $s(k)$.